OpenStudy (anonymous):

Points A and B lie on the graph of the equation y = −4.3x. What are the y-coordinates for each point? Point A: (−3, ) Point B: (5, )

OpenStudy (cookie_kvng):

@Willie579

OpenStudy (willie579):

What grade? I'm in 7th. :/

OpenStudy (willie579):

@June-27-2001

OpenStudy (anonymous):

i dont know how to do this and im in 9th grade

OpenStudy (cookie_kvng):

i'm in 9th too its easy sumtimes but the way u have it is confusing me :s

OpenStudy (malcolmmcswain):

This is really simple, all you have to do is solve for y by plugging in x. x = -3*-4.3 = 12.9

OpenStudy (alexandervonhumboldt2):

substitute the x coordinate into the equation: y=-4.3*(-3)=12.9 y=-4.3*5=-21.5

OpenStudy (anonymous):

thank you guys so much and whos in k12

OpenStudy (malcolmmcswain):

I used to be...

OpenStudy (anonymous):

i dont like it did you

OpenStudy (malcolmmcswain):

Not really XD It was a lot of work. I'm in high school now, and I do Plato and it's waaay harder. I wish I could do k12 again

OpenStudy (malcolmmcswain):

So, enjoy it while you can... school is hard.

OpenStudy (anonymous):

i dont like it and im in 9th grade and its hard